# STEPC: A Pixel-wise Nonuniformity Correction Framework for Photon-Counting CT in Multi-material Imaging Scenarios

### v1

@@ 3 Methods
where $S(E)$ is the input X-ray spectrum, $E$ denotes the photon energy, $R_{j}(E^{\prime},E)$ represents the pixel-specific photon-counting detector response function, $E^{\prime}$ denotes the detected energy, $E_{\max}$ is the maximum photon energy determined by the tube voltage. Assuming the existence of an ideal or reference response function $R_{j}^{ideal}(E^{\prime},E)$ that ensures uniformity across pixels, the corresponding ideal detector counts is:
[…]
Thus, the nonuniformity response error is:
[…]
We propose a STEPC method that directly utilizes multi-energy projection information to predict and correct detector nonuniformity for multi-material imaging. The detailed workflow is illustrated in Figure 2 and consists of the following steps:
Step 1: Ideal projection estimation. STEPC first measure the transmission through different combinations of multiple material slab phantoms. The log-transformed flat-field corrected projection data are then computed as $P(E_{k})=-\log(I_{E}^{FF})$ to reduce the influence of X-ray source and detector instabilities. Next, a second-order polynomial fit is applied to the 2D projection image $(x,y)$ for each energy bin $E_{k}$ to obtain the ideal projection $P_{ideal}(x,y,E_{k})$:
$P_{ideal}(x,y,E_{k})=\sum_{i+j\leq 2}b_{ij}^{(k)}x^{i}y^{j}$ (13)
Here, $(x,y)$ are the pixel coordinates, and the coefficients $b_{ij}^{(k)}$ are obtained by minimizing the mean square error between $P_{E_{k}}$ and $P_{ideal}(E_{k})$. We use MATLAB’s regress function [30] to solve for these coefficients. As shown in Figure 2(Step1), this polynomial fitting compensates for spatial variations in X-ray intensity and angular-dependent transmission thickness in the slab phantom, without requiring precise geometric measurements, thereby simplifying the calibration process.
$\Delta P_{ideal}(x,y,E_{k})=P(x,y,E_{k})-P_{ideal}(x,y,E_{k})$ (14)
As shown in Figure 2(Step2), the nonuniformity errors is dependent on the incident spectrum, which varies with material type and thickness. For single-material cases, single-energy projections can roughly estimate the error. However, for multi-materials, single-energy projections fail to capture the error accurately. In contrast, it illustrates that dual-energy information more effectively characterizes the error surface, allowing accurate modeling of errors caused by multi-materials.
Step 3: Empirical calibration model fitting. An empirical polynomial model is used to predict nonuniformity error. By incorporating all energy-channel projections as input, the model captures spectral variations associated with different materials, enabling more accurate and robust correction in multi-material imaging scenarios. The model is formulated as follows:
$\hat{\Delta P}(x,y,E_{k})=\sum_{|\boldsymbol{\alpha}|\leq p}c_{\boldsymbol{\alpha}}^{(k)}\prod_{n=1}^{N}P(x,y,E_{n})^{\alpha_{n}}$ (15)
where $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{N})$ is the multi-index vector, $|\boldsymbol{\alpha}|=\sum_{n=1}^{N}\alpha_{n}\leq p$ is the total polynomial degree, $N$ is the number of energy thresholds, and $c_{\boldsymbol{\alpha}}^{(k)}$ are the polynomial coefficients for energy channel $E_{k}$. These coefficients are determined by minimizing the mean square error of the estimator for ideal error $\Delta P_{ideal}$:
$MSE=\mathbb{E}\left[\left(\hat{\Delta P}(x,y,E_{k})-\Delta P_{ideal}(x,y,E_{k})\right)^{2}\right]$ (16)
Step 4: Correction. Finally, real object projection data are corrected by directly inputting the raw multi-energy projections into the fitted polynomial model to estimate the nonuniformity error, which is then subtracted to yield the corrected projection:
$P_{corr}(x,y,E_{k})=P(x,y,E_{k})-\hat{\Delta P}(x,y,E_{k})$ (17)
Notably, STEPC avoids the need for accurate thickness or density measurements of calibration phantoms by using 2D polynomial fitting to generate ideal reference projections. In addition, despite having only two energy thresholds, STEPC can incorporate iodine-based calibration to capture spectral variations from contrast agents, soft tissue, and bone, enabling effective correction in complex imaging scenarios.
Figure 3: Micro-PCCT system and multi-material slab phantoms. (a) Micro-PCCT imaging system; (b) slab phantom placement setup; (c) three types of slab phantoms: PMMA, aluminum, and iodixanol; (d) schematic of iodixanol phantom container; (e) combinations of PMMA and aluminum slabs with varying thicknesses; (f) combinations of iodixanol solution slabs with PMMA and aluminum slabs.
4 Experiments
4.1 Micro Photon Counting CT and Calibration Materials
As shown in Figure 3(a), the Micro Photon Counting CT (Micro-PCCT) system used in our experiments was jointly developed by Hainan University and United Imaging Life Science Instrument (LSI, Wuhan, China). It adopts a translate-rotate architecture where the object remains stationary and the gantrol rotates, reducing motion artifacts during imaging. The X-ray source was operated at 80 kV and 200 µA, and a 0.5 mm aluminum filter was used to suppress low-energy photons. The photon counting detector has a resolution of $2063\times 505$ pixels (after cropping peripheral invalid pixels), each with a 100 $\mu$$\times$ 100 $\mu$m pixel size, and supports two independently adjustable energy thresholds. The detector thresholds were set to 15 keV and 30 keV. Each energy channel uses a 12-bit counter capable of recording up to 4096 photons per acquisition. Each acquisition thus provides three energy bins: Total (15–80 keV) and High (30–80 keV), with the Low (15–30 keV) bin derived by subtraction: $\mathrm{Low}=\mathrm{Total}-\mathrm{High}$.
@@ 4.1 Micro Photon Counting CT and Calibration Materials
Three types of slab phantoms were designed, as shown in Figure 3(c), including PMMA slabs with thicknesses of 0, 5, 10, 15, 20, 30, and 40 mm; aluminum slabs with thicknesses of 0, 0.5, 1, 1.5, 2, 3, 4, 5, 6, and 8 mm; and iodixanol solution slabs with concentrations of 0, 5, 10, 15, 20, 30, 50, 70, 100, 150, and 250 mg/cm3, sealed in 2 mm thick PMMA containers with a 6 mm solution core (Figure 3(d)). Two imaging scenarios were considered. For the non-contrast scenario, calibration was performed using combinations of PMMA and aluminum slabs, as shown in Figure 3(e). For the iodine-enhanced scenario, the same PMMA and aluminum slab combinations were used, and additional calibration was performed by incorporating iodixanol solutions combined with PMMA and aluminum slabs (0,0, 0,1, 0,3, 10,0, 10,1, and 20,0), as illustrated in Figure 3(f). For each slab combination, 600 projection frames were acquired and averaged to reduce noise.
4.2 Phantom and Mouse Imaging
@@
To evaluate correction performance under different conditions, five cylindrical phantoms were scanned (Figure 4): PMMA-only cylinders, 200 mg/mL CaCl2 cylinders (PMMA container with 2 mm thickness), 50 mg/mL Iodixanol cylinder (PMMA container with 2 mm thickness), cylinders with inserts of CaCl2 at $\{100,200,400,600\}$ mg/mL, cylinders with inserts of CaCl2 $\{200,400\}$ mg/mL and iodixanol $\{20,50,100\}$ mg/mL. All phantoms had an outer diameter of 30 mm, and the inserts diameter is 8 mm. Projections were acquired in “Continuous” mode with 1440 views per rotation (0.25° per view), a field of view (FOV) of 50 mm, a source-to-isocenter distance (SID) of 90 mm, and a source-to-detector distance (SDD) of 325 mm. All other scan settings were the same as in the calibration step.
@@
4.3 Data Processing and Reconstruction
Detector nonuniformity correction was performed in the projection domain. The detector comprises $16\times 2$ tiles, separated by gap pixels and affected by defective pixels. These bad pixels were pre-identified and corrected via linear interpolation. Image reconstruction was carried out using the open-source TIGRE toolbox [31], employing the FDK algorithm with Hann windowing, resulting in volumes of size $1529\times 1529\times 400$. Images were converted to Hounsfield Units (HU) using a water phantom for calibration. All processing was performed in MATLAB R2024a [32].
@@ 4.3 Data Processing and Reconstruction
4.4 Baselines
We compare our STEPC method with four baselines: FF, STC, ATC, and PMDC. Since these methods work in different domains, appropriate post-processing was applied to ensure a fair comparison:
  • FF: Logarithmic projections were directly computed as $P_{E}^{FF}=-\log(I_{E}^{FF})$.
  • STC: Projections were computed using the calibrated thickness model: $P_{E}^{STC}=-\log\left(\frac{N_{E}^{STC}}{\bar{N}_{E,\mathrm{air}}}\right)$, where $N_{E}^{STC}=\bar{C}_{E}e^{\bar{A}_{E}T_{E}}$ and $\bar{C}_{E}$, $\bar{A}_{E}$ are average calibration coefficients obtained across all pixels.
  • ATC: Instead of using mean counts as described in Ref. [16], the same second-order 2D polynomial fit was applied to generate reference photon counts. After ATC correction in the counting domain, logarithmic air normalization was performed: $P_{E}^{ATC}=-\log\left(\frac{N_{E}^{ATC}}{\acute{N}_{E,\mathrm{air}}}\right)$. $\acute{N}_{E,\mathrm{air}}$ was also obtained by applying a second-order 2D polynomial fit to the air scan projection.
@@ 4.4 Baselines
    $P_{E(k)}=\sum_{i+j\leq 3}c_{ij}^{(k)}(T_{\mathrm{PMMA}})^{i}(T_{\mathrm{Al}})^{j},$ (18)
@@
    After the PMDC correction obtaining $T_{\mathrm{PMMA}}^{PMDC}$ and $T_{\mathrm{Al}}^{PMDC}$, averaged coefficients $\bar{c}_{ij}^{(k)}$ across all pixels were used to compute corrected projections:
    $P_{E(k)}^{PMDC}=\sum_{i+j\leq 3}\bar{c}_{ij}^{(k)}(T_{\mathrm{PMMA}}^{PMDC})^{i}(T_{\mathrm{Al}}^{PMDC})^{j}.$ (19)
Note: FF requires only air scans for calibration; STC used only PMMA slabs for calibration; ATC and STEPC used PMMA and aluminum slabs for the non-contrast scenario and additional iodixanol slabs for the contrast scenario; PMDC used only PMMA and aluminum for both scenarios due to its two-threshold limitation. The polynomial order for both PMDC and STEPC was set to three.
4.5 Metrics
To assess the proposed correction method, we calculated the Mean Local Standard Deviation (MLSD) using a $20\times 20$ pixels window across each corrected projections. Lower values indicate better projection uniformity after correction.
To assess ring artifact suppression in reconstructed images, we used the Ring Artifact Deviation (RAD) metric described by Rodesch et al. [11]. As shown in Figure 5, images were first converted to polar coordinates centered on the rotation axis. Angular averaging was then applied, followed by a second-order polynomial fit along the radial direction to remove cupping artifacts. Finally, the standard deviation of the residuals across valid radial ranges and all slices was calculated. Since the single material phantom and multi-material insert phantoms have different internal structures, we selected different ROI sizes when computing the RAD metric to ensure the regions uniformly, so that the measurement reflected only ring artifacts intensity. For the single-material cylindrical phantoms, a 20 mm diameter ROI was selected, while for the insert phantoms, an 8 mm ROI was used. For the mouse scans, due to the complexity of anatomical structures, smaller ROIs were chosen to ensure regional uniformity. Specifically, a 3.5 mm diameter ROI was used for the head scan, and a 3.7 mm diameter ROI was used for the kidney scan, as indicated by the red circles in Figure 11. However, the same ROI was applied across all methods for each dataset, ensuring fair and consistent comparisons.
5 Results
@@
We evaluated the corrected projection uniformity for different material slabs, and ring artifact suppression in cylindrical phantoms and mouse images. First, we quantitatively assessed projection uniformity for all PMMA+Al slab combinations (Figures 7) and PMMA+Al+iodixanol phantom slabs (Figures 8), with averaged MLSD results shown in Table 1. Second, we compared ring artifact suppression qualitatively (Figures 91011) and quantitatively (Tables 2) for phantoms and mouse imaging. Finally, we performed a sensitivity analysis (Table 3, Figure 12) to evaluate the impact of polynomial degree and calibration material choice on difference material objects imaging.
5.1 Projection Uniformity for Calibration Slabs

### v2

@@ IV-A Micro Photon Counting CT and Calibration Materials
Three types of slab phantoms were designed, as shown in Fig. 3(b), including PMMA slabs with thicknesses of 0, 5, 10, 15, 20, 30, and 40 mm; aluminum slabs with thicknesses of 0, 0.5, 1, 1.5, 2, 3, 4, 5, 6, and 8 mm; and iodixanol solution slabs with concentrations of 0, 5, 10, 15, 20, 30, 50, 70, 100, 150, and 250 mg/cm3, sealed in 2 mm thick PMMA containers with a 6 mm solution core (Fig. 3(c)). The slab phantoms were positioned approximately parallel to the detector plane, as illustrated in Fig. 3(a). Two imaging scenarios were considered. For the non-contrast scenario, calibration was performed using combinations of PMMA and aluminum slabs, as shown in Fig. 3(d). For the iodine-enhanced scenario, the same PMMA and aluminum slab combinations were used, and additional calibration was performed by incorporating iodixanol solutions combined with PMMA and aluminum slabs (0,0, 0,1, 0,3, 10,0, 10,1, and 20,0), as illustrated in Fig. 3(e). For each slab combination, 600 projection frames were acquired and averaged to reduce noise.
IV-B Phantom and Mouse Imaging
@@
To evaluate correction performance under different conditions, five cylindrical phantoms were scanned (Fig. 4): PMMA-only cylinders, 200 mg/mL CaCl2 cylinders (PMMA container with 2 mm thickness), 50 mg/mL Iodixanol cylinder (PMMA container with 2 mm thickness), cylinders with inserts of CaCl2 at $\{100,200,400,600\}$ mg/mL, cylinders with inserts of CaCl2 $\{200,400\}$ mg/mL and iodixanol $\{20,50,100\}$ mg/mL. All phantoms had an outer diameter of 30 mm, and the inserts diameter is 8 mm. Projections were acquired in “Continuous” mode with 1440 views per rotation (0.25° per view), a field of view (FOV) of 50 mm, a source-to-isocenter distance (SID) of 90 mm, and a source-to-detector distance (SDD) of 325 mm. All other scan settings were the same as in the calibration step.
@@
IV-C Data Processing and Reconstruction
Detector nonuniformity correction was performed in the projection domain. The detector comprises $16\times 2$ tiles, separated by gap pixels and affected by defective pixels. These gap and defective pixels were pre-identified and corrected via linear interpolation. Image reconstruction was carried out using the open-source TIGRE toolbox [5], employing the FDK algorithm with Hann window filtering. The reconstructed image matrix size was $1529\times 1529\times 400$. The reconstruction FOV was $50~\mathrm{mm}\times 50~\mathrm{mm}\times 10~\mathrm{mm}$ for the cylindrical phantoms and $35~\mathrm{mm}\times 35~\mathrm{mm}\times 8~\mathrm{mm}$ for the mice. Images were converted to Hounsfield Units (HU) using a water phantom for calibration. All processing was performed in MATLAB R2024a.
@@ IV-C Data Processing and Reconstruction
IV-D Baselines
We compared our STEPC method with four baselines: FF, STC, ATC, and PMDC. Since these methods work in different domains, appropriate post-processing was applied to ensure a fair comparison:
  • FF: Logarithmic projections were directly computed as $P_{E}^{\mathrm{FF}}=-\log(I_{E}^{\mathrm{FF}})$.
  • STC: Projections were computed using the calibrated thickness model: $P_{E}^{\mathrm{STC}}=-\log\left(\frac{N_{E}^{\mathrm{STC}}}{\bar{N}_{E,\mathrm{air}}}\right)$, where $N_{E}^{\mathrm{STC}}=\bar{C}_{E}e^{\bar{A}_{E}T_{E}}$ and $\bar{C}_{E}$, $\bar{A}_{E}$ are average calibration coefficients obtained across all pixels.
  • ATC: Instead of using mean counts as described in Ref. [29], the same second-order 2D polynomial fit was applied to generate reference photon counts. After ATC correction in the counting domain, logarithmic air normalization was performed: $P_{E}^{\mathrm{ATC}}=-\log\left(\frac{N_{E}^{\mathrm{ATC}}}{\acute{N}_{E,\mathrm{air}}}\right)$. $\acute{N}_{E,\mathrm{air}}$ was also obtained by applying a second-order 2D polynomial fit to the air scan projection.
@@ IV-D Baselines
    $P_{E(k)}=\sum_{1\leq i+j\leq p}c_{ij}^{(k)}(T_{\mathrm{PMMA}})^{i}(T_{\mathrm{Al}})^{j},$ (25)
@@
    After the PMDC correction obtaining $T_{\mathrm{PMMA}}^{\mathrm{PMDC}}$ and $T_{\mathrm{Al}}^{\mathrm{PMDC}}$, averaged coefficients $\bar{c}_{ij}^{(k)}$ across all pixels were used to compute corrected projections:
    $P_{E(k)}^{\mathrm{PMDC}}=\sum_{1\leq i+j\leq p}\bar{c}_{ij}^{(k)}(T_{\mathrm{PMMA}}^{\mathrm{PMDC}})^{i}(T_{\mathrm{Al}}^{\mathrm{PMDC}})^{j}.$ (26)
    When $p=1$, the model yields an equivalent monoenergetic projection, enabling beam hardening correction. To ensure a fair comparison with other methods that do not perform beam hardening correction, $p=3$ was used to represent the non beam hardening corrected case.
Note: FF requires only air scans for calibration. STC uses only PMMA slabs, as PMMA closely approximates the attenuation of soft biological tissue and is commonly used for single-material calibration. ATC and STEPC used PMMA and aluminum slabs for the non-contrast scenario and additional iodixanol slabs for the contrast scenario; PMDC, limited by its two-threshold constraint, uses PMMA and aluminum slabs for both scenarios, with PMMA and aluminum serving as proxies for soft and high-density tissues, respectively, which is the most common choice. This limitation is inherent to PMDC and restricts its applicability to iodine-enhanced imaging, rather than a design choice in our experimental comparison.
IV-E Metrics
To assess the proposed correction method, two complementary image quality metrics were employed: the Mean Local Standard Deviation (MLSD) for projection-domain uniformity assessment, and the Ring Artifact Deviation (RAD) for reconstruction-domain artifact quantification.
(1) Mean Local Standard Deviation (MLSD). For each corrected projection image, projection uniformity is commonly evaluated using variance or standard deviation metrics, as reported in Ref. [29]. However, to properly account for variations in slab phantom thickness across different geometric angles, we adopt the MLSD metric, which quantifies local intensity consistency. Specifically, a sliding window of $20\times 20$ pixels was applied across the projection image, and the standard deviation $\sigma_{n}$ within each window $n$ was calculated. The MLSD was then defined as the mean of all local standard deviations:
${\rm MLSD}=\frac{1}{N}\sum_{n=1}^{N}\sigma_{n},$ (27)
where $N$ denotes the total number of valid local windows. A lower MLSD value indicates higher spatial uniformity of the projection, reflecting more effective correction of detector nonuniformity.
(2) Ring Artifact Deviation (RAD). To assess ring artifact suppression in reconstructed images, the Ring Artifact Deviation (RAD) metric provides an effective quantitative measure of ring artifact severity in central uniform regions, as described by Rodesch et al. [31]. As shown in Fig. 5, each reconstructed slice $I_{\text{real}}(s,x,y)$ was first transformed from Cartesian coordinates $(x,y)$ to polar coordinates $(r,\theta)$ centered on the rotation axis, yielding $I_{\text{polar,real}}(s,r,\theta)$. The angular mean intensity was then computed as:
$\bar{I}_{\text{polar,real}}(s,r)=\frac{1}{N_{\theta}}\sum_{\theta=1}^{N_{\theta}}I_{\text{polar,real}}(s,r,\theta),$ (28)
where $N_{\theta}$ is the total number of angular samples. A second-order polynomial was subsequently fitted along the radial direction to approximate the ideal, ring-free intensity distribution $\bar{I}_{\text{polar,ideal}}(s,r)$. The RAD metric was then defined as the standard deviation of the residual intensity differences across all valid radii $r$ and slices $s$:
${\rm RAD}={\rm STD}_{r,s}\!\left\{\bar{I}_{\text{polar,real}}(s,r)-\bar{I}_{\text{polar,ideal}}(s,r)\right\}.$ (29)
Lower RAD values correspond to more effective suppression of ring artifacts. Since the single-material and multi-material phantoms possess different internal structures, distinct circular regions of interest (ROIs) with radius $r$ were selected to ensure uniform sampling and fair comparison. For single-material cylindrical phantoms, a diameter of 20 mm was used; for insert phantoms, a diameter of 8 mm was chosen. For in-vivo mouse scans, smaller ROIs were defined due to anatomical heterogeneity: 3.5 mm for the head region and 3.7 mm for the kidney region, as indicated by the red circles in Fig. 12. The same ROI radius was applied across all compared methods within each dataset to maintain consistency and fairness.
V Results
@@
We evaluated the corrected projection uniformity for different material slabs, and ring artifact suppression in cylindrical phantoms and mouse images. First, simulation experiments were conducted to assess the fitting accuracy of the second-order 2D polynomial model for estimating ideal projections (Fig. 6, and Table I). Projection uniformity was then quantitatively evaluated for all PMMA+Al slab combinations (Fig. 7, 8) and PMMA+Al+iodixanol phantom slabs (Fig. 9), with averaged MLSD results shown in Table II. Second, we compared ring artifact suppression qualitatively (Fig. 101112) and quantitatively (Tables III) for phantoms and mouse imaging. In addition, we demonstrated the potential of the proposed method for beam hardening correction (Fig. 13) and evaluated its impact on material decomposition accuracy (Fig. 14 and Table IV). Finally, we performed a sensitivity analysis (Fig. 15, and Table V) to evaluate the impact of polynomial degree and calibration material choice on different material objects imaging.
V-A Ideal Projection Fitting Accuracy Evaluation
Figure 6: Second-order 2D polynomial fitting results for simulated projection data of a slab phantom (PMMA: 20 mm, Al: 3 mm) under a $0^{\circ}$ tilt angle.
TABLE I: Relative errors of second-order 2D polynomial fitting for simulated projection data.
tilt angles Combinations RE($\times 10^{-4}$) (MEAN[MIN,MAX])
Low High Total
$0^{\circ}$ PMMA+AL 0.82[0.77,0.86] 0.81[0.79,0.83] 0.80[0.77,0.84]
PMMA+AL+I 0.82[0.79,0.84] 0.81[0.79,0.85] 0.82[0.78,0.84]
$2.5^{\circ}$ PMMA+AL 0.82[0.76,0.90] 0.89[0.87,0.91] 0.83[0.77,0.90]
PMMA+AL+I 0.85[0.77,0.89] 0.89[0.85,0.91] 0.85[0.78,0.89]
$5^{\circ}$ PMMA+AL 1.47[1.08,2.00] 1.96[1.78,2.09] 1.64[1.32,2.02]
PMMA+AL+I 1.65[1.18,1.92] 1.92[1.54,2.08] 1.73[1.36,1.96]
The proposed STEPC framework employs a second-order 2D polynomial surface to represent the ideal projection. To evaluate the fitting accuracy of this model under different material and thickness combinations of slab phantoms, as well as under varying tilt angles, a series of simulation experiments were conducted.
The X-ray source spectrum was generated using the Spektr software [30] with a tube voltage of 80 kV and a 0.5 mm Al filter. The energy response function of the photon-counting detector was adopted from Ref. [33], with two energy thresholds set at 15 keV and 30 keV. The mass attenuation coefficients of PMMA, Al, and I were obtained from the NIST database [15]. The SID and SDD were set to 140 mm and 325 mm, respectively. All thickness and material combinations were identical to those described in Section IV-A. Three different phantom tilt angles were investigated ($0^{\circ}$, $2.5^{\circ}$, and $5^{\circ}$). Since the primary objective of this experiment was to evaluate the projection fitting accuracy of the second-order 2D polynomial surface, noise effects were not considered. The projection fitting accuracy was quantitatively assessed by computing the relative error (RE) between the fitted ($P_{\mathrm{fit}}$) and simulated projection ($P_{\mathrm{simu}}$) values for different material and thickness combinations:
$\mathrm{RE}_{s}=\frac{1}{N_{xy}}\sum_{x,y}\frac{\left|P_{\mathrm{fit}}(x,y,s)-P_{\mathrm{simu}}(x,y,s)\right|}{P_{\mathrm{simu}}(x,y,s)},$ (30)
Here, $(x,y)$ are the pixel coordinates, $N_{xy}$ is the total number of pixels, and $s$ indexes the calibration combinations of multiple material slabs. As an illustrative example, the fitted and simulated 2D projection images for a slab phantom combination with a PMMA thickness of 20 mm and an Al thickness of 3 mm at a tilt angle of $0^{\circ}$ are shown in Fig. 6, demonstrating excellent agreement between the two projections. The quantitative results are summarized in Table I. When no phantom tilt was introduced, the average RE across all energy projections was approximately $0.81\times 10^{-4}$. This value increased only marginally to $0.85\times 10^{-4}$ at a tilt angle of $2.5^{\circ}$, and remained below $2.09\times 10^{-4}$ even at $5^{\circ}$. These results demonstrate that the proposed second-order 2D polynomial surface can accurately estimate ideal projections for slab phantoms with different materials and thicknesses, even in the presence of moderate geometric tilts. Importantly, this approach does not require precise phantom positioning or detailed geometric measurements, thereby significantly simplifying the calibration process.
V-B Projection Uniformity for Calibration Slabs